OpenStudy (anonymous):

(x+4)²/12+(y-1)²/8=1 What the the coordinates of the center? What is the length of the major & minor axis?

OpenStudy (anonymous):

center is \((-4,1)\)

OpenStudy (anonymous):

perfect, thank you

OpenStudy (anonymous):

you can read it off from the equation

OpenStudy (anonymous):

\(a^2=12\) so \(a=\sqrt{12}\) and the length of the major axis is \(2\sqrt{12}\)

OpenStudy (anonymous):

oh i see how to find it now

OpenStudy (anonymous):

but it is different for the minor axis i suppose

OpenStudy (anonymous):

@satellite73